five distinct letters are to be transmitted through a communication channel.a total number of 15 blanks is to be inserted between the 2 letters with atleast 3 between every 2.the number of ways in which this can be done is? please reply soon.............

Grade:

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear Sri Valli,

Ans:- The min no of blanks between every two letter are=3 and as the total no of blanks are 15 and hence the max no of blanks that may contain there are=6. Hence using Binomial and GP series expansion we get the no of ways to place the 15 gaps are = 6C3= 20 and the no of ways of arrangeing 5 different letters are= 5!=120

Hence the no of different arrangements are=(120×20)=2400(ans)

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best  Sri Valli!!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur